OpenStudy (anonymous):

Solve the system of equations and choose the correct ordered pair. 6x - 6y + 11 = 17 9y = 3x + 15 A. (6, 5) B. (4, 3) C. (3, 2) D. (6, 3)

OpenStudy (anonymous):

Well the first thing would be to simplify the first equation, and then solve for x; 6x - 6y + 11 = 17 6x - 6y = 17 - 11 6x - 6y = 6 6x = 6y + 6 6y + 6 x = ------- 6 x = y + 1 Now we enter it in the other equation; 9y = 3x + 15 9y = 3(y + 1) + 15 9y = 3y + 3 + 15 9y - 3y = 18 6y = 18 y = 18/6 y = 3 x = y + 1 x = 3 + 1 x = 4 The answer would be that x = 4 and y = 3, so it be (4,3) which is answer choice b